OpenStudy (anonymous):

Which of the following could be the ratio of the length of the longer leg of a 30-60-90 triangle to the length of its hypotenuse? Check all that apply.

OpenStudy (anonymous):

\[1: \sqrt{3}\] \[3\sqrt{3}:6\] \[\sqrt{3}:\sqrt{3}\] \[\sqrt{2}:\sqrt{3}\] \[3:2\sqrt{3}\] \[\sqrt{3}:2\]

OpenStudy (anonymous):

rad3 to 2

OpenStudy (anonymous):

look at the 30-60-90 triangle